Please help me solve this question:
How many mm of a 5% HCl mixture and a 20% HCl mixture must be combined to get a 10 milimeters of a 12.5% HCl mixture?

Answer by stanbon(75887)

How many mm of a 5% HCl mixture and a 20% HCl mixture must be combined to get a 10 milimeters of a 12.5% HCl mixture?
------
Equation:
acid + acid = acid
0.05x + 0.20(10-x) = 0.125*10
----
50x + 200*10 - 200x = 125*10
----
-150x = -75*10
x = (1/2)10
x = 5 ml (amt of 5% mixture needed.
10-x = 5 ml (amt. of 20% mixture needed).
============
Cheers,
Stan H.
=============